If the sound intensity is 6.8 x 10-10 W/m2, and assuming area of eardrum is 0.9 cm2, the energy falling on the human eardrum per second is (write your answer in terms of 10-14 joules):
Answer #1

Given that -10 1 = 6.810 W/m² * = 0,9*10-4 m2 AE AL -10 IA = 6.8x0,9 x 10 X10 6,8x0,97 74 ΛΕ = 6.12 x 1014 w -14 Ils ΔΕ ΔΕ 6,